# Quantum entanglement in phase space

### v2

@@ Methods
Reference [12] presents a more general entanglement criterion based on reduced states. The author demonstrates that negativity of the reduced Wigner function in mode $a_{+}$ implies partial transpose negativity. This approach extends previous work [11] by incorporating additional local unitaries before and after passing the two modes through a balanced beam-splitter, corresponding to symplectic transformations of the coordinate system. Wigner negativity in one of the beam-splitter output modes indicates entanglement. This criterion is again a special case of our lower bound, as we consider both unitaries and antiunitaries. Additionally, we provide upper bounds to the reduced Wigner function, further complementing the study of the relationship between this method based on the Wigner function and partial transposition.
VII VII. Comparison with Bell inequality and EPR steering criterion based on Wigner functions
@@
VII.1 A. EPR criterion based on pseudospin measurements
In [37], it is proposed the EPR steering criterion in terms of pseudospin measurements
@@
$W_{\Pi^{x}}(x,p)$ $=-\delta(x)\mathcal{P}\frac{1}{p},$ (46)
@@
VII.2 B. CHSH-type Bell inequality based on Wigner fucntions
@@
VIII VIII. Detailed analysis of some exemplary quantum states
VIII.1 A. two-mode squeezed thermal states
We start with two-mode squeezed thermal (TMST) states to investigate the applicability of our criteria to Gaussian states. An arbitrary TMST states can be generated by a specific scheme in [37, 38]. A single-mode phase-insensitive Gaussian channel is applied to one of the modes of a two-mode squeezed vacuum state with squeezing parameter $s$. The channel is decomposed into a quantum-limited attenuator with transmissivity $\eta$ followed by a quantum limited amplifier with gain $\cosh^{2}r$, introducing loss and noise to the states, respectively.
@@
The Wigner functions of Gaussian states can be writtern as functions of the covariance matrix
@@
$W_{AB}(s,\eta,r)$ $=\frac{1}{4\pi^{2}}\frac{e^{-\frac{-\left(\left(x_{B}^{2}+p_{B}^{2}\right)\left(\eta+(\eta-2)\cosh{(2r)}\right)\right)+\left(2\left(x_{A}^{2}+p_{A}^{2}\right)+\left(x_{B}^{2}+p_{B}^{2}\right)\eta+\left(x_{B}^{2}+p_{B}^{2}\right)\eta\cosh{(2r)}\right)\cosh{(2s)}+4(p_{A}p_{B}-x_{A}x_{B})\sqrt{\eta}\cosh{(r)}\sinh{(2s)}}{2\left(2\eta\cosh^{2}{(r)}-(\eta+(-2+\eta)\cosh{(2r)})\cosh{(2s)}\right)}}}{\left(\cosh^{2}{(r)}\left(\eta-(\eta-1)\cosh{(2s)}\right)+\cosh{(2s)}\sinh^{2}{(r)}\right)}$ (52)
@@
which indicates that the entanglement of TMSV states is detected if $s>0$.
@@
Although useful for some states, criterion (II) no more provides a tight bound for all TMST states. Under the optimal linear transformation $(x^{\prime}=x,p^{\prime}=-p^{\prime})$, the optimal parameter $\theta_{\text{opt}}$ in criterion (II) can be expressed as a function of $(s,\eta,r)$
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$
$=\frac{1}{16\pi^{2}}e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)}\left[(1+\epsilon)\left(x_{A}^{2}x_{B}^{2}+x_{A}^{2}p_{B}^{2}+p_{A}^{2}x_{B}^{2}+p_{A}^{2}p_{B}^{2}\right)+4\epsilon(x_{A}x_{B}-p_{A}p_{B})-2\epsilon\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)+4\epsilon\right].$ (57)
@@
$\iint_{-\infty}^{\infty}W_{AB}\left(\frac{\sqrt{2}}{2}x,\frac{\sqrt{2}}{2}p,\frac{\sqrt{2}}{2}x,-\frac{\sqrt{2}}{2}p\right)=\iint_{-\infty}^{\infty}W_{AB}\left|{\left(\frac{\sqrt{2}}{2}x,\frac{\sqrt{2}}{2}p,\frac{\sqrt{2}}{2}x,-\frac{\sqrt{2}}{2}p\right)}\right|=\frac{1+3\epsilon}{4\pi}\leq\frac{1}{2\pi}.$ (58)
Both criteria (I,II) can reveal entanglement for $\rho(\Phi^{+})$ for $\epsilon>1/3$. This bound is equivalent with the known tight bound obtained by PPT criterion [39].
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$
$=\frac{1}{16\pi^{2}}e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)}\left[(1-\epsilon)\left(x_{A}^{2}x_{B}^{2}+x_{A}^{2}p_{B}^{2}+p_{A}^{2}x_{B}^{2}+p_{A}^{2}p_{B}^{2}\right)+4\epsilon(x_{A}x_{B}+p_{A}p_{B})+2\epsilon\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)-4\epsilon\right].$ (59)
@@
In conclusion, our criteria, which perform as well as PPT criterion, are necessary and sufficient conditions for the considered Werner states. In addition, we have analyzed some known entanglement criteria for better comparison in Fig. 8. It’s found that Duan-Simon criteria [1, 2] are unable to detect entanglement for any $\epsilon$, and the entanglement based on quantum Fisher information [40] and Hillery-Zubairy criteria [50] are useful but not tight (the former reveals entanglement for both $\rho(\Phi^{+})$ and $\rho(\Psi^{+})$ when $\epsilon\gtrsim 0.65$, and the latter can only detect entanglement for $\rho(\Psi^{+})$ when $\epsilon>\frac{\sqrt{5}-1}{2}$ ).
VIII.3 C. partially dephased two-mode cat states
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$ $=\frac{e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}+8\gamma^{2}\right)}}{4\pi^{2}\left(e^{4\gamma^{2}}+1\right)}\left(e^{8\gamma^{2}}\epsilon\cos{(2\gamma(p_{A}+p_{B}))}+\left(e^{4\gamma^{2}}-\epsilon+1\right)\cosh{(2\gamma(x_{A}+x_{B}))}\right).$ (62)
@@
which indicates the violation for any $\gamma>0$ and $\epsilon\in[0,1]$. Criterion (II) will yield the same results if an integration over the entire space is considered. We conclude that both criteria (I,II) can provide tight bounds for $\rho(\psi^{+})$ (see also Fig. 3 in the main text).
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})=\frac{e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}+8\gamma^{2}\right)}}{4\pi^{2}\left(e^{4\gamma^{2}}-1\right)}\left(-e^{8\gamma^{2}}\epsilon\cos{(2\gamma(p_{A}+p_{B}))}+\left(e^{4\gamma^{2}}+\epsilon-1\right)\cosh{(2\gamma(x_{A}+x_{B}))}\right).$ (64)
Under the linear transformation $(x^{\prime}=-x,p^{\prime}=-p)$, the Wigner function is negative at $\epsilon=1$ (see Fig. 7(a)), we thus calculate criterion (III)
$\iint_{-\infty}^{\infty}W_{AB}\left(x,p,-x,-p\right)\text{d}x\text{d}p=\frac{e^{-4\gamma^{2}}\left(1-\left(1+e^{4\gamma^{2}}\right)\epsilon\right)}{4\pi}\geq 0.$ (65)
The entanglement is revealed only if $\epsilon>1/(1+e^{4\gamma^{2}})$. We also calculate criterion (II) numerically, under the transformation $(x^{\prime}=x,p^{\prime}=p)$ and $\theta=\pi/4$. As illustrated in Fig. 7(b) , the minimum $\epsilon_{\min}$ for violating criterion (II) coincides with the one for violating criterion (III).
In Fig. 8, as a summary, we compare the entanglement detection capabilities of our criteria (I,II,III) to Duan-Simon [1, 2], Hillery-Zubairy [50], QFI [40], and PPT [14, 39] criteria, respectively. For each of the considered states, at least one criterion out of the family (I,II,III) can provide a tight bound. Note that, different from QFI criterion [40] and PPT criterion [14, 39] which require full information of the density matrix, the criteria (I,II,III) only needs partial information determined by the transformation $(x^{\prime},p^{\prime})$.
IX IX. Error analysis
IX.1 A. Discretization
@@
In an practical scenario, we imagine that the experiment can sample $W_{AB}$ on a square grid of spacing $\delta$ over the finite region $R$. Specifically, we define the discrete points along the $x$-axis as $x_{j}=x_{\text{min}}+j\delta$ with $j=0,1,2,\cdots,(x_{\text{max}}-x_{\text{min}})/\delta$, and analogously, the discrete points $p_{k}$ along the $p$-axis, yielding the discrete Wigner functions $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$. In Fig. 9, we compare the discrete Wigner functions $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$ under sampling resolution $\delta=0.2$ and $\delta=0.4$, showing that the Wigner functions become coarser as $\delta$ grows.
@@
Fig. 10 illustrates the amount of violation $I_{D}$ and relative error $\epsilon=|I_{D}-I|/|I|$ as a function of the sampling resolution $\delta$. It is found that our criterion is robust against finite sampling under realistic discrete approximations, with the relative error as small as $\epsilon\approx 2\times 10^{-3}$ even for a coarse sampling $\delta=0.5$.
We observe from Fig. 9 that the Wigner function $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$ for pure cat states $|\psi\rangle$ remains non-negative over the entire phase space. Similar non-negative distribution can also be found in $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$ for two-mode squeezed vacuum (TMSV) states and Bell states $|\Phi^{+}\rangle$. In both these cases, we restrict the sampling region $R$ to $x,p\in[-3,3]$, and observe the violation values increase slightly as $\delta$ grows. At $\delta=0.5$, the relative error is found to be $\epsilon\approx 3\times 10^{-4}$ for the TMSV states and $\epsilon\approx 6\times 10^{-4}$ for Bell states $|\Phi^{+}\rangle$.
On the other hand, for the Bell state $|\Psi^{+}\rangle$ the joint Wigner function $W_{AB}(x,p,x,p)$ in criterion (II) involves both positive and negative values. Fig. 11(a) shows the discrete Wigner functions $W_{AB}(x_{j},p_{k},x_{j},p_{k})$ for a sampling resolution of $\delta=0.2$. Fig. 11(b) presents how the amount of violation changes with sampling resolution $\delta$. Contrary to the non-negative case, the change of violation is no longer monotonic with the sampling resolution $\delta$, reflecting the contribution of both positive and negative pixels. Within sampling resolution $\delta\in[0,0.5]$, we find the largest relative error $\epsilon\approx 0.04$ at $\delta=0.4$. For $\delta<0.12$, the observed violation clearly approaches the ideal value given by $\delta\rightarrow 0$ to very good approximation.
@@
To illustrate our approach, we consider the two-mode squeezed vacuum (TMSV) state with Wigner function $W(\alpha;\beta)=\frac{4}{\pi^{2}}\exp\left[-2\cosh 2r\left(|\alpha|^{2}+|\beta|^{2}\right)+2\sinh 2r\left(\alpha\beta+\alpha^{*}\beta^{*}\right)\right]$. We evaluate the CHSH combination $\mathcal{B}$ as
@@
According to Bell’s theorem, classical correlations satisfly $2\leq\mathcal{B}\leq 2$. For the TMSV state with fixed squeezing parameter $r$, the optimal displacement amplitude $\mathcal{J}_{\mathrm{opt}}(r)=\frac{2r-\ln(\cosh 2r)}{3e^{2r}-e^{-2r}}$ maximizes the violation of the CHSH inequality [8, 9]. This choice yields
@@
We only need to make sure that the displacements are sampled uniformly over the desired region. These measurements are statistically equivalent to evaluating the Wigner function at a single point. We now show that this randomized measurement scheme significantly reduces the required number of measurements. From Eqs. (69,75), the variance is

### v3

@@ VI. A comparison between criterion (III) and previous criteria based on Wigner functions
Reference [12] presents a more general entanglement criterion based on reduced states. The author demonstrates that negativity of the reduced Wigner function in mode $a_{+}$ implies partial transpose negativity. This approach extends previous work [11] by incorporating additional local unitaries before and after passing the two modes through a balanced beam-splitter, corresponding to symplectic transformations of the coordinate system. Wigner negativity in one of the beam-splitter output modes indicates entanglement. This criterion is mathematically equivalent to our lower bound. Additionally, we provide upper bounds to the reduced Wigner function, further complementing the study of the relationship between this method based on the Wigner function and partial transposition.
VII. Comparison with Bell inequality and EPR steering criterion based on Wigner functions
@@
A. EPR criterion based on pseudospin measurements
In [50, 38], it is proposed the EPR steering criterion in terms of pseudospin measurements
@@ A. EPR criterion based on pseudospin measurements
$W_{\Pi^{y}}(x,p)$ $=-\delta(x)\mathcal{P}\frac{1}{p},$ (49)
@@
B. CHSH-type Bell inequality based on Wigner functions
@@
VIII. Detailed analysis of some exemplary quantum states
A. two-mode squeezed thermal states
We start with two-mode squeezed thermal (TMST) states to investigate the applicability of our criteria to Gaussian states. An arbitrary TMST state can be generated by a specific scheme in [38, 39]. A single-mode phase-insensitive Gaussian channel is applied to one of the modes of a two-mode squeezed vacuum state with squeezing parameter $s$. The channel is decomposed into a quantum-limited attenuator with transmissivity $\eta$ followed by a quantum limited amplifier with gain $\cosh^{2}r$, introducing loss and noise to the states, respectively.
@@ VIII. Detailed analysis of some exemplary quantum states
The Wigner functions of Gaussian states can be written as functions of the covariance matrix
@@
$W_{AB}(s,\eta,r)=\frac{1}{4\pi^{2}}\frac{1}{\left(\cosh^{2}{(r)}\left(\eta-(\eta-1)\cosh{(2s)}\right)+\cosh{(2s)}\sinh^{2}{(r)}\right)}\times$
$e^{-\frac{-\left(\left(x_{B}^{2}+p_{B}^{2}\right)\left(\eta+(\eta-2)\cosh{(2r)}\right)\right)+\left(2\left(x_{A}^{2}+p_{A}^{2}\right)+\left(x_{B}^{2}+p_{B}^{2}\right)\eta+\left(x_{B}^{2}+p_{B}^{2}\right)\eta\cosh{(2r)}\right)\cosh{(2s)}+4(p_{A}p_{B}-x_{A}x_{B})\sqrt{\eta}\cosh{(r)}\sinh{(2s)}}{2\left(2\eta\cosh^{2}{(r)}-(\eta+(-2+\eta)\cosh{(2r)})\cosh{(2s)}\right)}}$ (55)
@@
whose violation for $s>0$ indicates that the entanglement of TMSV states is detected.
@@
Although useful for some states, criterion (II) no more provides a tight bound for all TMST states. Under the optimal linear transformation $(x^{\prime}=x,p^{\prime}=-p)$, the optimal parameter $\theta_{\text{opt}}$ in criterion (II) can be expressed as a function of $(s,\eta,r)$
@@ B. Werner states
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$ $=\frac{1}{16\pi^{2}}e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)}\Big[(1+\epsilon)\left(x_{A}^{2}x_{B}^{2}+x_{A}^{2}p_{B}^{2}+p_{A}^{2}x_{B}^{2}+p_{A}^{2}p_{B}^{2}\right)$
$\quad+4\epsilon(x_{A}x_{B}-p_{A}p_{B})-2\epsilon\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)+4\epsilon\Big].$ (60)
@@
$\iint_{-\infty}^{\infty}W_{AB}\left(\frac{\sqrt{2}}{2}x,\frac{\sqrt{2}}{2}p,\frac{\sqrt{2}}{2}x,-\frac{\sqrt{2}}{2}p\right)\mathrm{d}x\mathrm{d}p$
$=\iint_{-\infty}^{\infty}\left|{W_{AB}\left(\frac{\sqrt{2}}{2}x,\frac{\sqrt{2}}{2}p,\frac{\sqrt{2}}{2}x,-\frac{\sqrt{2}}{2}p\right)}\right|\mathrm{d}x\mathrm{d}p$
$=\frac{1+3\epsilon}{4\pi}$
$\leq\frac{1}{2\pi}.$ (61)
Both criteria (I,II) are violated for $\epsilon>1/3$, revealing entanglement in $\rho(\Phi^{+})$. This bound is equivalent with the known tight bound obtained by PPT criterion [40].
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$ $=\frac{1}{16\pi^{2}}e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)}\Big[(1-\epsilon)\left(x_{A}^{2}x_{B}^{2}+x_{A}^{2}p_{B}^{2}+p_{A}^{2}x_{B}^{2}+p_{A}^{2}p_{B}^{2}\right)$
$\quad+4\epsilon(x_{A}x_{B}+p_{A}p_{B})+2\epsilon\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}\right)-4\epsilon\Big].$ (62)
@@
In conclusion, our criteria, which perform as well as PPT criterion, are necessary and sufficient conditions for the considered Werner states. In addition, we have analyzed some known entanglement criteria for better comparison in Fig. 4 of the main text. It’s found that Duan-Simon criteria [1, 2] are unable to detect entanglement for any $\epsilon$, and the entanglement based on quantum Fisher information [41] and Hillery-Zubairy criteria [42] are useful but not tight (the former reveals entanglement for both $\rho(\Phi^{+})$ and $\rho(\Psi^{+})$ when $\epsilon\gtrsim 0.65$, and the latter can only detect entanglement for $\rho(\Psi^{+})$ when $\epsilon>\frac{\sqrt{5}-1}{2}$ ).
C. partially dephased two-mode cat states
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$ $=\frac{e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}+8\gamma^{2}\right)}}{4\pi^{2}\left(e^{4\gamma^{2}}+1\right)}$
$\times\left(e^{8\gamma^{2}}\epsilon\cos{(2\gamma(p_{A}+p_{B}))}+\left(e^{4\gamma^{2}}-\epsilon+1\right)\cosh{(2\gamma(x_{A}+x_{B}))}\right).$ (65)
@@
which is violated for any $\gamma>0$ and $\epsilon\in(0,1]$. Criterion (II) will yield the same results if an integration over the entire space is considered. We conclude that both criteria (I,II) can provide tight bounds for $\rho(\psi^{+})$ (see also Fig. 3 in the main text).
@@
$W_{AB}(x_{A},p_{A},x_{B},p_{B})$ $=\frac{e^{-\frac{1}{2}\left(x_{A}^{2}+p_{A}^{2}+x_{B}^{2}+p_{B}^{2}+8\gamma^{2}\right)}}{4\pi^{2}\left(e^{4\gamma^{2}}-1\right)}$
$\times\left(-e^{8\gamma^{2}}\epsilon\cos{(2\gamma(p_{A}+p_{B}))}+\left(e^{4\gamma^{2}}+\epsilon-1\right)\cosh{(2\gamma(x_{A}+x_{B}))}\right).$ (67)
Since the Wigner function under the linear transformation $(x^{\prime}=-x,p^{\prime}=-p)$ is negative at $\epsilon=1$ (see Fig. 8(a)), we calculate criterion (III)
$\iint_{-\infty}^{\infty}W_{AB}\left(x,p,-x,-p\right)\text{d}x\text{d}p=\frac{e^{-4\gamma^{2}}\left(1-\left(1+e^{4\gamma^{2}}\right)\epsilon\right)}{4\pi}.$ (68)
For separable states this quantity must be non-negative, so entanglement is revealed when $\epsilon>1/(1+e^{4\gamma^{2}})$. We also calculate criterion (II) numerically, under the transformation $(x^{\prime}=x,p^{\prime}=p)$ and $\theta=\pi/4$. As illustrated in Fig. 8(b) , the minimum $\epsilon_{\min}$ for violating criterion (II) coincides with the one for violating criterion (III).
In Fig. 4 of the main text, as a summary, we compare the entanglement detection capabilities of our criteria (I,II,III) to Duan-Simon [1, 2], Hillery-Zubairy [42], QFI [41], and PPT [23, 40] criteria, respectively. For each of the considered states, at least one criterion out of the family (I,II,III) can provide a tight bound. Note that, different from QFI criterion [41] and PPT criterion [23, 40] which require full information of the density matrix, the criteria (I,II,III) only needs partial information determined by the transformation $(x^{\prime},p^{\prime})$.
IX. Error analysis
A. Discretization
@@
In a practical scenario, we imagine that the experiment can sample $W_{AB}$ on a square grid of spacing $\delta$ over the finite region $R$. Specifically, we define the discrete points along the $x$-axis as $x_{j}=x_{\text{min}}+j\delta$ with $j=0,1,2,\cdots,(x_{\text{max}}-x_{\text{min}})/\delta$, and analogously, the discrete points $p_{k}$ along the $p$-axis, yielding the discrete Wigner functions $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$. In Fig. 9, we compare the discrete Wigner functions $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$ under sampling resolution $\delta=0.2$ and $\delta=0.4$, showing that the Wigner functions become coarser as $\delta$ grows.
@@
Fig. 10 illustrates the amount of violation $I_{D}$ and relative error $\epsilon_{\text{rel}}=|I_{D}-I|/|I|$ as a function of the sampling resolution $\delta$. It is found that our criterion is robust against finite sampling under realistic discrete approximations, with the relative error as small as $\epsilon_{\text{rel}}\approx 2\times 10^{-3}$ even for a coarse sampling $\delta=0.5$.
We observe from Fig. 9 that the Wigner function $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$ for pure cat states $|\psi\rangle$ remains non-negative over the entire phase space. Similar non-negative distribution can also be found in $W_{AB}(x_{j},p_{k},x_{j},-p_{k})$ for two-mode squeezed vacuum (TMSV) states and Bell states $|\Phi^{+}\rangle$. In both these cases, we restrict the sampling region $R$ to $x,p\in[-3,3]$, and observe the violation values increase slightly as $\delta$ grows. At $\delta=0.5$, the relative error is found to be $\epsilon_{\text{rel}}\approx 3\times 10^{-4}$ for the TMSV states and $\epsilon_{\text{rel}}\approx 6\times 10^{-4}$ for Bell states $|\Phi^{+}\rangle$.
On the other hand, for the Bell state $|\Psi^{+}\rangle$ the joint Wigner function $W_{AB}(x,p,x,p)$ in criterion (II) involves both positive and negative values. Fig. 11(a) shows the discrete Wigner functions $W_{AB}(x_{j},p_{k},x_{j},p_{k})$ for a sampling resolution of $\delta=0.2$. Fig. 11(b) presents how the amount of violation changes with sampling resolution $\delta$. Contrary to the non-negative case, the change of violation is no longer monotonic with the sampling resolution $\delta$, reflecting the contribution of both positive and negative pixels. Within sampling resolution $\delta\in[0,0.5]$, we find the largest relative error $\epsilon_{\text{rel}}\approx 0.04$ at $\delta=0.4$. For $\delta<0.12$, the observed violation clearly approaches the ideal value given by $\delta\rightarrow 0$ to very good approximation.
@@ B. How to measure the criteria
To illustrate our approach, we consider the two-mode squeezed vacuum (TMSV) state with Wigner function $W(\alpha;\beta)=\frac{4}{\pi^{2}}\exp\left[-2\cosh 2r\left(|\alpha|^{2}+|\beta|^{2}\right)+2\sinh 2r\left(\alpha\beta+\alpha^{*}\beta^{*}\right)\right]$ where $r$ is the squeezing parameter. We evaluate the CHSH combination $\mathcal{B}$ as
@@
According to Bell’s theorem, classical correlations satisfy $2\leq\mathcal{B}\leq 2$. For the TMSV state with fixed squeezing parameter $r$, the optimal displacement amplitude $\mathcal{J}_{\mathrm{opt}}(r)=\frac{2r-\ln(\cosh 2r)}{3e^{2r}-e^{-2r}}$ maximizes the violation of the CHSH inequality [8, 9]. This choice yields
@@
We only need to make sure that the displacements are sampled uniformly over the desired region. These measurements are statistically equivalent to evaluating the Wigner function at a single point. We now show that this randomized measurement scheme significantly reduces the required number of measurements. From Eqs. (72,78), the variance is
@@
Another notable distinction from CHSH inequality tests is that while the CHSH violation has an upper bound of $2\sqrt{2}$, the violation of our entanglement criteria has no such upper limit. For states with squeezing approaching infinity, the violation of our criteria can grow unboundedly. So the entanglement detection becomes more robust to errors: if the experiment is limited by some systematic or statistical errors in Wigner integration, one can compensate by generating larger squeezing to yield a larger violation value over the separable threshold.
@@
C. Additional comments on error analysis and number of measurements
Figure 12: Illustration of (a) standard detection and (b) randomized detection. In (a) standard detection, the phase space is discretized into $M\sim D^{2}$ samples, and $N_{\text{shot}}$ repeated shots are required for each sample. In (b) randomized detection, $N_{\text{shot}}$ random shots are performed within the chosen space (yellow solid circle). Criteria (I) and (III) require the average values over all statistics, which is equivalent to $M=1$ expectation value. Criterion (II) requires the average values over positive and negative regions, respectively, corresponding to $M=2$ expectation values. In both methods, the amount of total data is $M\cdot N_{\text{shot}}$.
In the preceding two subsections, we analyzed errors that are independent of specific experimental apparatus details but may still play a dominant role in practical implementations. We now consider the overall uncertainty including apparatus-level contributions. We characterize the combined uncertainty from random and systematic errors as $\sigma_{\mathrm{tot}}=\sqrt{\sigma_{\mathrm{stat}}^{2}+\sigma_{\mathrm{syst}}^{2}}$. According to the discussion in the previous subsections, if grid discretization is employed to evaluate the two-dimensional integral, the associated error can be classified as systematic, since the bias originates from the finite grid spacing, i.e., the discretization approximation due to non-infinitesimal $\delta$. In contrast, if randomized sampling measurements are used to estimate the integral over the integration region, the error can be classified as statistical, as it arises from finite repetitions and decreases with increasing sample size. Beyond these considerations, the experimental apparatus itself may introduce additional systematic errors, such as parity readout errors and displacement calibration errors.
We now elaborate on how to compare measurement counts. In typical direct Wigner function measurement experiments, the total number of measurements can be written as $N_{\mathrm{tot}}=M\cdot N_{\mathrm{shot}}$, where $M$ is the number of sampled displacement points and $N_{\mathrm{shot}}$ is the number of repeated shots at each displacement point. For full tomography of a two-mode quantum state, if the effective truncation dimension of each mode is $D$, the required number of measurement settings $M$ typically scales as $D^{4}$. In contrast, under our grid discretization scheme from the previous subsections, we effectively reconstruct a two-dimensional phase-space distribution, so $M$ scales as $D^{2}$. The advantage of this approach is that it provides the most straightforward way to measure and compute the various criteria when no prior knowledge of the quantum state’s principal distribution region in phase space is available. However, when the experimentalist knows in advance which quantum state will be prepared and has prior information about the principal phase-space distribution region, employing randomized measurements can substantially reduce the required $M$.
Specifically, in the randomized measurement protocol, the number of measurement rounds can be taken as $M=1$ for criteria (I) and (III), and $M=2$ for criterion (II). This is because we employ randomized sampling: within a single measurement round, $N_{\mathrm{shot}}$ randomly displaced sampling points are included, with each sampling point measured only once (yielding a single $\pm 1$ outcome) rather than repeating $N_{\mathrm{shot}}$ times at the same point. Thus, one measurement round still corresponds to $N_{\mathrm{shot}}$ shots. Criteria (I) and (III) involve linear integrals, and the required data is essentially a single mean value over the two-dimensional phase space. Estimating this mean via randomized sampling is statistically equivalent to directly measuring the Wigner function at a single point. Criterion (II), on the other hand, requires partitioning the distribution into positive and negative regions based on prior knowledge, and performing mean-value measurements separately over each region, thus necessitating two measurement rounds. Even if the prior partitioning is imprecise or the sign of certain regions is misjudged, this will only lower the computed value of criterion (II) rather than increase it, thereby weakening the detection capability without falsely indicating entanglement.